A cage is lowered down a pit shaft, 600m deep in 46 seconds. What is its average speed in meters per second and kilometers per hour?
1
Expert's answer
2019-12-16T11:11:19-0500

Average speed:


"v = \frac{600 \, m}{46 \, s}= 13 \frac{m}{s} =13 \frac{\frac{km}{1000}}{\frac{hour}{3600}} = 47 \frac{km}{hour}"